OpenStudy (anonymous):

i need help!.... Write the equation of the line which passes through (4, –2) and is perpendicular to the line with equation 2y – x = 1.

jimthompson5910 (jim_thompson5910):

First solve for y \[\Large 2y – x = 1\] \[\Large 2y = x + 1\] \[\Large y = \frac{x + 1}{2}\] \[\Large y = \frac{1}{2}x + \frac{1}{2}\] What is the slope in that last equation above?

OpenStudy (anonymous):

x + 2y = 2 ??

jimthompson5910 (jim_thompson5910):

no, that's not the final answer

OpenStudy (anonymous):

2x - y = 4 ???

jimthompson5910 (jim_thompson5910):

still no

OpenStudy (anonymous):

x - 2y =6

jimthompson5910 (jim_thompson5910):

that line is parallel and not perpendicular to the given one

OpenStudy (anonymous):

so 2x + y = 6

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

any way of tellin if this is true or false?

jimthompson5910 (jim_thompson5910):

What is this asking or making a statement about?

jimthompson5910 (jim_thompson5910):

I just see a table, but no statement. So what is it saying true or false about?

OpenStudy (anonymous):

it only asks if the table is t or f ..??

jimthompson5910 (jim_thompson5910):

is this the whole question?

OpenStudy (anonymous):

sorry the statement above says, "the table below does not dfine a funciton" t or f ??

jimthompson5910 (jim_thompson5910):

what do you think

jimthompson5910 (jim_thompson5910):

are there any cases where an x value maps to more than one y value?

OpenStudy (anonymous):

i said true

jimthompson5910 (jim_thompson5910):

why don't you think it's a function?

OpenStudy (anonymous):

because x values match to more than one y??

OpenStudy (anonymous):

or some x's do

jimthompson5910 (jim_thompson5910):

are there any cases where one x value maps to more than one y value?

OpenStudy (anonymous):

yes, to the 1

jimthompson5910 (jim_thompson5910):

no, you're looking at it backwards from what I'm asking

OpenStudy (anonymous):

ohhhh sorry, that is the range.. then it would be no

jimthompson5910 (jim_thompson5910):

each x maps to only one y value

jimthompson5910 (jim_thompson5910):

you don't have x = 1 for example mapping to two different things (say y = -1 and y = 0). It only maps to one thing and that is y = -1

OpenStudy (anonymous):

so the table is true, it is a function

jimthompson5910 (jim_thompson5910):

It is a function

OpenStudy (anonymous):

or excuse me, false because it it is a function

jimthompson5910 (jim_thompson5910):

The original statement was "the table below does not define a function"

jimthompson5910 (jim_thompson5910):

So the original statement is false

OpenStudy (anonymous):

right, thanks!

jimthompson5910 (jim_thompson5910):

yw